OpenStudy (anonymous):

At t = 0, a particle leaves the origin with a velocity of 9.0 m/s in the positive y direction and moves in the xy plane with a constant acceleration of (2.0i  4.0j) m/s2. At the instant the x coordinate of the particle is 15 m, what is the speed of the particle?

OpenStudy (anonymous):

Use the formula s=at^2/2 <=> sqrt(2s/a)=t So know you know the time. You will need to use the basic v=a*t+Vo formula, and learn horizontal and verticall speed( Y speed, and X speed if you prefer ), and once you know them you add them using Pitagoras Law ( meaning you take the square root of the sum their second powers to know the actuall speed of the particle) Hope i helped .